OpenStudy (anonymous):

I really need some help! I already did the work and all but, I need someone to check it because I think I got it wrong!

OpenStudy (anonymous):

Here is the question Each tennis ball is 2.63 inches in diameter. A sketch of the specially designed container is below. How many tennis balls can fit inside the container? How many more tennis balls could fit into the container if the container’s dimensions are doubled? You must show all work to receive credit. Here is my work: Cylinder: Height 36 in/ Diameter 13 in 3.14 (7.5²) (36) 3.14 (56.25) (36) 176.714586764*36 Volume=6361.72512352 Cone: Height 8 in/ Diameter 13 in V=1/3 (3.14r²)(h) 1/3(3.14*7.5²) (8) 1/3(3.14*56.25) (8) 1/3 (176.714586764) (8) 1/3 (1413.716694112) Volume=471.2388980373333333 A+B= 6832.9640215573333333 Tennis Ball: Diameter 2.63 V=4/3 3.14r³ 4/3 3.14*1.315³ 4/3 3.14*2.273930875 4/3* 7.1401429475 V=9.5201905966666667 6832.9/9.5= 717 tennis balls If you double it you would be able to fit 717 more tennis balls in the container

zepdrix (zepdrix):

Hmm I don't understand what the `cone` part is used for.

OpenStudy (anonymous):

It's a cylinder with a cone on top of it.

zepdrix (zepdrix):